OpenStudy (anonymous):

Find the product. (3a – b)3

OpenStudy (anonymous):

Is the 3 the power, of are you multiplying it by 3?

OpenStudy (anonymous):

is it (3z - b) * 3 ?

OpenStudy (anonymous):

@tgawade @larryboxaplenty 3 is an exponent

OpenStudy (anonymous):

or you can write it as (3a -b)^3 = (3a - b)*(3a- b)*(3a - b) & then use FOIL method ,,

OpenStudy (anonymous):

(x - y)^3 = x^3 - y^3 - 3xy (x - y)

OpenStudy (anonymous):

just replace x by 3a & y by b